OpenStudy (clamin):

You are MIguel Cervantes de Navas y Colon, captain in the Royal Spanish Army in Sevilla in the year 1842. Outside your barracks window, there is s a stack of cannnonballs, as shown . On an idle afternoon you decide to calculated how many cannonballs are in the stack.. What is the number of cnnonballs? a. 347 b.361 c.285 d.385

Directrix (directrix):

http://openstudy.com/study#/updates/537db367e4b0e90e4b2dcc85 may help, especially for the diagram. I counted 10 layers in the stack. Each stack contains a perfect square of balls. 10^2 + 9^2 + .... + 1^2 = (10 * 11 * 21) /6 = ?

Directrix (directrix):

P(n) = n(n + 1)(2n + 1)/6 is the formula for the sum of the first n perfect squares. It can be proved by mathematical induction. @clamin